# Artificial Intelligence LLM Large Language Model

### v2

@@ Appendix E Exploratory Analyses
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
[…]
@@
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
Table E2. Welch’s independent t-test comparisons (between-subjects) of the customized LLMaaS chatbot (UniGPT) and ChatGPT Free, with participants using only one system ($n_{\text{UniGPT}}=70$, $n_{\text{ChatGPT Free}}=176$)

### v3

@@ Appendix F Exploratory Analyses
Notes. 1 Holm-adjusted significance threshold per hypothesis rank (sequential correction). 2 Cohen’s d is reported as the standardized effect size for each comparison. 3 Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons. 4 Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test. 5 Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
Table F2. Welch’s independent t-test comparisons (between-subjects) of the customized LLMaaS chatbot (UniGPT) and ChatGPT Free, with participants using only one system ($n_{\text{UniGPT}}=70$, $n_{\text{ChatGPT Free}}=176$)
@@
Notes. 1 Holm-adjusted significance threshold per hypothesis rank (sequential correction). 2 Cohen’s d is reported as the standardized effect size for each comparison. 3 Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons. 4 Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test. 5 Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.